A Bibliography of Publications in Biometrics: 1940–1949

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

25 June 2020 Version 1.00

Title word cross-reference

p [12, 8]. p^{3} [12]. p^{8} [9]. p^{n} [8, 9].

additivity [15]. analysis [2, 6, 12, 11]. assumptions [2, 6].

between [13]. binomial [13]. block [5]. blocks [12, 8, 9]. both [14].

chiasma [7]. Comparing [11]. comparisons [4]. consequences [2].

degree [15]. designs [5, 12, 8, 9, 9].

error [14].

Fitting [14]. formation [7]. freedom [15].

REFERENCES 2

general [12, 8, 9]. genetic [7]. II [9]. III [12]. incomplete [5]. India [5]. individual [11, 4]. Introduction [8]. lattice [12, 8, 9]. lattices [1, 10]. line [14]. logarithmic [13]. means [11]. methods [4]. negative [13]. non [15]. non-additivity [15]. One [15]. plots [12, 8, 9]. Poisson [13]. power [12, 8, 9]. Preliminary [1]. prime [12, 8, 9]. prime-power [12, 8, 9]. Probability [4]. quantitative [7]. ranking [4]. recombination [7]. rectangular [1, 10]. relation [13]. replicates [12]. report [1]. satisfied [2]. series [13]. Some [2]. squares [9]. straight [14]. subject [14]. tables [4]. theory [12, 7, 8, 9]. transformations [3]. Triple [10]. underlying [6]. use [3]. variables [14]. variance [2, 6, 11]. varieties [12, 8, 9]. work [5].

References

Harshbarger:1946:PRR

[1] Boyd Harshbarger. Preliminary report on the rectangular lattices. Biometrics, 2(??):115–119, 1946. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cochran:1947:SCW

[2] W. G. Cochran. Some consequences when the assumptions for the analysis of variance are not satisfied. Biometrics, 3(??):22–38, 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bartlett:1947:UT

[3] M. S. Bartlett. The use of transformations. Biometrics, 3(??):39–52, 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wilcoxon:1947:PTI

[4] Frank Wilcoxon. Probability tables for individual comparisons by ranking methods. Biometrics, 3(??):119–122, 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bose:1947:RWI

[5] R. C. Bose. Recent work on “incomplete block designs” in India. Biometrics, 3(??):176–178, 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Eisenhart:1947:AUA

[6] Churchill Eisenhart. The assumptions underlying the analysis of variance. Biometrics, 3(1):1–21, 1947. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fisher:1948:QTG

[7] R. A. Fisher. A quantitative theory of genetic recombination and chiasma formation. Biometrics, 4(??):1–13, ???? 1948. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kempthorne:1948:GTPa

[8] Oscar Kempthorne and Walter T. Federer. The general theory of primepower lattice designs. I. Introduction and designs for p^{n} varieties in blocks of p plots. Biometrics, 4(??):54–79, 1948. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kempthorne:1948:GTPb

[9] O. Kempthorne and W. T. Federer. The general theory of prime-power lattice designs. II. Designs for p^{n} varieties in blocks of p^{8} plots, and in squares. Biometrics, 4(??):109–121, 1948. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Harshbarger:1949:TRL

[10] Boyd Harshbarger. Triple rectangular lattices. Biometrics, 5(??):1–13, 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tukey:1949:CIM

[11] John W. Tukey. Comparing individual means in the analysis of variance. Biometrics, 5(??):99–114, 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Federer:1949:GTP

[12] Walter T. Federer. The general theory of prime-power lattice designs. III. The analysis for p^{3} varieties in blocks of p plots with more than 3 replicates. Biometrics, 5(??):144–161, 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Quenouille:1949:RBL

[13] M. H. Quenouille. A relation between the logarithmic, Poisson, and negative binomial series. Biometrics, 5(??):162–164, 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bartlett:1949:FSL

[14] M. S. Bartlett. Fitting a straight line when both variables are subject to error. Biometrics, 5(??):207–212, 1949. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tukey:1949:ODF

[15] John W. Tukey. One degree of freedom for non-additivity. Biometrics, 5(??):232–242, ???? 1949. CODEN BIOMB6. ISSN 0006-341X (print),

1541-0420 (electronic). URL http://www.jstor.org/stable/3001938.